Mr. Reyes estimates he can do a job in 12 hours while Mr. Tan estimates he can do the same job in 15 hours. How long would it take them to finish the job if they work together?
Answer by edjones(7311)

Reyes does 1/12 job/hr
Tan does 1/15 job/hr
together 1/12 + 1/15 or 5/60+4/60=9/60=3/20 of the job in an hour.
t=time in hrs for entire job.
(3/20)*t
t=20/3=6 2/3 hrs
Ed